A local high school needs to hire several cafeteria workers and bus drivers. Cafeteria workers earn R13 800


per month and bus drivers earn R17 250 per month. The school board gave the school permission to spend


no more than R345 000 on the salaries per month, but must hire more than 15 people.


Let B represent the number of bus drivers hired.


Let C represent the number of cafeteria workers hired.


For every bus driver hired there must be at least two cafeteria workers hired. The number of cafeteria


workers must at most be three more than twice the number of bus drivers.


Write a system of linear inequalities that the school could use to determine the number of cafeteria workers


and bus drivers they can hire

1
Expert's answer
2022-04-20T02:46:00-0400

"17250b+13800c\le 345000\\
b+c\gt 15 \\
c\ge 2b \\
c\le 2b+3 \\"


One can check that the following pairs are solutions: b=5 and c=11; b=5 and c=12; b=5 and c=13; b=6 and c=12; b=6 and c=13; b=6 and c=14; b=6 and c=15; b=7 and c=14; b=7 and c=15; b=7 and c=16.